ok last one for the day. Thank you for your time.
type an integer or simplified fraction x=?
-8x-21=-23 2/3

Answer by nycsharkman(136)

Solve for x: -8x-21=-23 2/3
The first thing you want to do is rewrite the mixed number -23_2/3 as an improper fraction.
So, -23_2/3 becomes -(71/3).
We now have -8x - 21 = -(71/3)
We multiply every term on BOTH sides of the fractional equation by -3 to remove the fraction on the right side.
It is easier working without fractions.
-8x times -3 = 24x
-21 times - 3 = 63
-(71/3) times - 3 = 71
We now have this linear equation:
24x + 63 = 71
Subtract 63 from 71 and then equate difference to 24x.
24x = 8
Divide both sides by 24 to find x.
x = 8/24
x = 1/3
Got it?
-